OpenStudy (anonymous):

the perimeter of a rectangle equals the sum of the lengths of its four side. if the width of rectangle is 8 inches and its perimeter is 40 inches, find the length of the rectangle.

OpenStudy (anonymous):

im lost

OpenStudy (mathstudent55):

Let the length be L Let the width be W Let the perimeter be P P = L + L + W + W Right?

OpenStudy (anonymous):

P=2L+2W

OpenStudy (mathstudent55):

That's good. Now substitute all variables you know by the numbers you are given: W = 8 and P = 40

OpenStudy (mathstudent55):

40 = 2L + 2(8) Now solve for L

OpenStudy (anonymous):

16

OpenStudy (mathstudent55):

\(16 + 16 + 8 + 8 \ne 40 \)

OpenStudy (anonymous):

im lost

OpenStudy (mathstudent55):

\(40 = 2L + 2(8)\) \(40 = 2L + 16\) \(24 = 2L\) \(12 = L\) The length is 12 inches.

OpenStudy (anonymous):

8

OpenStudy (anonymous):

im lost

OpenStudy (primeralph):

https://maps.google.com/

OpenStudy (anonymous):

primeralph what you want me to do

OpenStudy (primeralph):